OpenStudy (camerondoherty):

The lengths of two sides of a triangle are shown below: Side 1: \(5x^2 + 2x - 3\) Side 2: \(8x - 3x^2 + 2\) The perimeter of the triangle is \(2x^3 - 5x^2 + 4x - 8\) What is the total length of the two sides, 1 and 2, of the triangle?

OpenStudy (nincompoop):

that's easy

OpenStudy (camerondoherty):

Um, I was thinking \[2x^2+10x-1\] Pretty sure i screwed up in my math though

OpenStudy (nincompoop):

do you think the perimeter has anything that you need to solve for the length of the two sides given?

OpenStudy (camerondoherty):

No, but there are a few other parts in the question tht may be of use in tht

OpenStudy (nincompoop):

your math is okay

OpenStudy (camerondoherty):

Oh fr? So I have it right? Wow, Thanks! :3

OpenStudy (camerondoherty):

If I wanted to find the missing angle would i subtract?

OpenStudy (camerondoherty):

Subtract the perimeter from the sum

OpenStudy (nincompoop):

angle?

OpenStudy (nincompoop):

yeah subtract

OpenStudy (camerondoherty):

*side

OpenStudy (nincompoop):

you mean side

OpenStudy (camerondoherty):

Yea Haha cx

OpenStudy (nincompoop):

side 1 + side 2 + side 3 = perimeter

OpenStudy (nincompoop):

perimeter - (side 1 + side 2) = side 3

OpenStudy (camerondoherty):

So... \[(2x^3 - 5x^2 + 4x - 8)-(2x^2+10x-1)\]

OpenStudy (nincompoop):

let us say A, B, C if A+B+C = P then A - A + B - B + C = P -A -B C = P-A -B

OpenStudy (nincompoop):

yes

OpenStudy (camerondoherty):

A-A B-B? Lmao brb

OpenStudy (nincompoop):

ye, because we want to isolate C by itself